in the right triangle the ratio of the measures of the three sides is 10:24:26, and its perimeter is 180 cm find the measure of the shortest side of the triangle
Answer by w_parminder(53)


One important thing about this question is that it can also be solved if it is not given that the triangle is a right triangle
Solution
Sides are in the ratio 10:24:26 so let the sides of the triangle are
10x, 24x & 26x
perimeter of triangle is 180 cm,
so 10x + 24x + 26x = 180
60x = 180
so x = 180/60 = 3
so the smaller side of the triangle i.e. 10x = 10*3
= 30 cm